OpenStudy (anonymous):

Please HELP! Medals included! Find each percent of change. Round to the nearest tenth of a percent. State if it is increase or decrease. From 109 to 98.

OpenStudy (anonymous):

change = 109-98 = 11 decrease percentage change = 11/109 x 100 = 10.091 = 10 % decrease.

OpenStudy (anonymous):

THANK YOU SO MUCH